OpenStudy (anonymous):

A garden table and a bench cost $1000 combined. The cost of the garden table is three times the cost of the bench. What is the cost of the bench?

OpenStudy (anonymous):

are there answer choices

OpenStudy (anonymous):

no theres not

OpenStudy (anonymous):

let me try to solve it for u

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Table= 3/4 (three quarters) Bench= 1/4 (One quarter) Divide the 1000 by 4 to get , = 250 To get the price of the table multiply by three (750) The bench is 250 Hope that helps! :)

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

@hype.child we are here to guide the asker to the correct answer. We are not here to just give them the answer. This is against the code of conduct. For further details refer to http://openstudy.com/code-of-conduct Do not commit such errors again. Thanks.

OpenStudy (anonymous):

u welcome

OpenStudy (anonymous):

Wow